Circular No. 7014

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1946+274 = GRO J1944+26 D. A. Smith, Massachusetts Institute of Technology (MIT); and T.

Takeshima, Goddard Space Flight Center (GSFC), NASA, on behalf of the RXTE/ASM and PCA teams at MIT and GSFC, communicate: “The ASM on RXTE has detected emission from a location that overlaps the previous localization of the 1976 Ariel V transient 3A 1942+274 (Warwick et al. 1981, MNRAS 197, 865; galactic latitude +1^{o}.7). The corners of the ASM 6′ × 26′ error region are at α = 19^{h}47^{m}.1, δ = +27^{o}16′; 19^{h}46^{m}.2, +27^{o}25′; 19^{h}45^{m}.5, +27^{o}25′; 19^{h}46^{m}.2, +27^{o}18′ (equinox 2000.0; 90-percent-confidence total error). The source first came to our attention at ∼ 90 mCrab (2–12 keV) on Sept. 15.91 UT, with a very hard spectrum. A retroactive search reveals that the 1-day average intensity has been rising from 13 mCrab on Sept. 5 to 60 mCrab on Sept. 15. The RXTE/PCA observed this location on Sept. 16.02 and recorded an intensity of ∼ 110 mCrab (2–60 keV) with a very strong 15.83-s (± 0.02 s) pulsation. The pulse profile is double-peaked, with a pulse fraction of ∼ 29 percent. There is, to our knowledge, no previous report of pulsations from 3A 1942+274. If this is a reoccurrence of the Ariel source, the 8′ × 52′ Ariel position restricts the location to the westernmost 5′ of the ASM region (see

http://xte.mit.edu/~dasmith/x1946 map.gif). Further PCA observa-

tions are in progress. Observations at other wavelengths are encouraged.”

C. A. Wilson, M. H. Finger, R. B. Wilson, and D. M. Scott, Marshall Space Flight Center, NASA, report for the BATSE team: “Pulsations from a previously-uncatalogued x-ray pulsar in the Vul-Cyg region have been detected with BATSE in the energy band 20–50 keV from Sept. 8 through Sept. 14. A preliminary location (95-percent confidence) is bounded by α = 19^{h}34^{m} to 19^{h}54^{m} and δ = +22^{o}.5 to +30^{o}.5 (equinox 2000.0). GRO J1944+26 rose from a 20–50-keV rms pulsed flux of < 15 mCrab on Sept. 7 to 27 ± 3 mCrab on Sept. 13. Observations between Sept. 8 and 14 are consistent with a barycentric pulse frequency of 0.06315331(6) Hz on Sept. 11.0 UT, with a frequency derivative of 2.2(2) × 10^{−12} Hz/s. The pulse profile in the band 20–50 keV has a single asymmetric peak.”

COMET C/1998 P1 (WILLIAMS) Visual m_{1} estimates: Aug. 26.38 UT, 7.6 (P. Williams, Heathcote,

N.S.W., 20×80 binoculars); Sept. 6.40, 7.4 (J. Shanklin, Hobart, Tasmania, 10×50 binoculars); 15.40, 8.2 (D. A. J. Seargent, The Entrance, N.S.W., 25×100 binoculars).

1998 September 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green